# BOTS: A Unified Framework for Bayesian Online Task Selection in LLM Reinforcement Finetuning

### v2

@@ 4.2 Fusing Two sources of Evidence: Analyzing the Impact of $\rho$
(2) Over-reliance on implicit evidence degrades long-term performance. As training progresses, settings with a large $\rho$ (e.g., 0.5, 1.0) show a declining ETR, as the implicit evidence cannot perfectly track the learning model’s fine-grained progress. It often fails to recognize when tasks are fully mastered ($p=1$), leading to their wasteful selection. This inaccuracy ultimately harms late-stage performance, as seen in the aggregated TTB and BSF metrics where large $\rho$ values underperform.
@@
4.3 Forgetting and Adaptivity: Analyzing the Impact of $\lambda$
Recall as defined in Equation 1, $\lambda$ controls the discounting of historical information in the Beta posterior. A smaller $\lambda$ places more emphasis on recent evaluations, improving adaptability but increasing the variance of the estimates. Conversely, a larger $\lambda$ gives greater weight to accumulated history, yielding more stable but less responsive estimates. Furthermore, Proposition 3.4 shows that $\lambda$ determines the scale of the equivalent counts in the Beta posterior. Intuitively, reducing $\lambda$ accelerates the decay of equivalent counts, increasing the randomness of posterior sampling, whereas enlarging $\lambda$ slows the decay and stabilizes the estimates.
@@ 4.3 Forgetting and Adaptivity: Analyzing the Impact of $\lambda$
We investigate the role of $\lambda$ in task selection by varying $\lambda\in\{0.0,0.05,0.1,0.2,0.5,1.0\}$, while keeping other hyperparameters fixed at their default values ($\rho=0.1$, posterior sampling enabled).
@@
Figure 3: Qwen2.5-1.5B-Instruct on Math. Ratio of sampled training tasks (measured over 16 rollouts) with passing rates: strictly between 0 and 1 (L1), strictly greater than 0 (L2), and strictly less than 1 (L3), along with MATH500 Accuracy (avg@1), plotted against training steps.
Benchmark MATH500 AIME24 Math Aggregated Performance
$\lambda=0.0,\rho=0.1$ 0.88 0.57 0.62 1.09 1.11 1.03 0.77 - - 1.25 1.50 0.75 0.87 0.58 0.71 1.11 1.08 1.04
$\mathbf{\lambda=0.05,\rho=0.1}$ 0.92 0.45 0.75 1.12 1.07 1.02 0.64 1.11 1.16 1.25 1.75 1.08 0.94 0.50 0.76 1.15 1.07 1.05
$\mathbf{\lambda=0.1,\rho=0.1}$ 0.89 0.49 0.57 1.13 1.05 1.05 0.51 1.00 1.00 1.25 1.75 1.00 0.89 0.56 0.64 1.12 1.07 1.05
$\mathbf{\lambda=0.2,\rho=0.1}$ 0.92 0.56 0.59 1.08 1.08 1.06 0.63 - - 0.75 2.00 0.75 0.95 0.58 0.66 1.07 1.11 1.06
$\lambda=0.5,\rho=0.1$ 1.11 0.53 0.77 1.11 1.07 1.03 0.51 0.89 - 1.25 1.75 0.83 1.14 0.74 0.86 1.12 1.06 1.03
$\lambda=1.0,\rho=0.1$ 1.15 0.68 0.80 1.04 1.04 1.03 0.60 0.63 - 1.00 2.50 0.83 1.14 0.68 1.00 1.04 1.09 1.02
@@
Table 2: TTB and BSF evaluated on MATH500, AIME24, and aggregated performance with Qwen2.5-1.5B-Instruct. For TTB, notation “-” indicates that the target performance is never achieved within the evaluation window. The best and second best results are marked accordingly.
@@
Analysis and Takeaways. The hyperparameter $\lambda$ governs the framework’s adaptivity to the non-stationary learning process. Our key takeaway is that task selection must be adaptive, and moderate forgetting is essential for long-term performance. An overly long memory ($\lambda\to 0$) prevents the system from recognizing when a task has been mastered, while an overly short memory ($\lambda\to 1$) leads to unstable estimates. This conclusion is supported by the following observations from Figure 3 and Table 2:
@@
(1) Small $\lambda$ struggles with mastered tasks: Settings with small $\lambda$ (e.g., 0.0, 0.05) exhibit a declining ETR in the mid-to-late training phase. This is because their long memory makes them slow to update the difficulty of tasks that the model has recently mastered (from $p<1$ to $p=1$), causing them to be wastefully selected. This ultimately limits late-stage performance gains.
(2) Large $\lambda$ lacks stability: Conversely, large $\lambda$ values (e.g., 0.5, 1.0) correspond to rapid forgetting. This makes the posterior estimates highly volatile and overly dependent on the most recent (and potentially noisy) observations. As a result, they fail to build stable beliefs about task difficulties, leading to only marginal improvements over the random baseline in the early training stages.
@@
(3) Moderate $\lambda$ balances stability and adaptivity: A moderate $\lambda\in\{0.05,0.1,0.2\}$ strikes a balance. It provides enough memory to form stable difficulty estimates while being adaptive enough to track the model’s evolving capabilities. This leads to strong and consistent performance gains across the entire training process, achieving the best TTB and BSF results.
@@ 4.4 Performance Comparison Across Models and Domains
Offline 0.96 0.51 0.78 1.16 1.12 1.01 0.68 1.03 - 1.09 1.00 0.99 1.23 1.36 - 0.67 0.78 0.89
BOTS-MoPPS 1.78 1.32 - 0.90 0.90 0.99 0.78 0.96 1.09 1.09 1.02 1.02 0.87 1.04 - 1.13 0.96 0.96
BOTS-DOTS 0.83 0.75 0.81 1.11 1.09 1.08 0.67 0.65 0.79 1.17 1.09 1.02 0.66 1.32 - 1.28 0.92 0.93
BOTS 0.89 0.56 0.64 1.12 1.07 1.05 0.58 0.90 0.77 1.17 1.08 1.03 0.85 0.94 1.05 1.19 1.01 1.00
@@
We conduct extended experiments across model sizes (1.5B and 7B) and task domains (math, code, logic). We compare BOTS under the default setting ($\lambda=0.1,\rho=0.1$, posterior sampling enabled) against four baselines. Two are out-of-framework baselines: (i) Random, where tasks are uniformly sampled, and (ii) Offline, where tasks are ranked from easy to hard based on success probabilities of Qwen2.5-7B-Instruct and Qwen3-30B-A3B (for tie-breaking). Two are within-framework baselines: (iii) BOTS-MoPPS, with $\lambda=0.0,\rho=0.0$ and posterior sampling enabled, which reduces our framework to MoPPS (qu2025can) and thus enables direct evaluation of a purely explicit-evidence strategy; and (iv) BOTS-DOTS, with $\lambda=1.0,\rho=1.0$ and posterior sampling disabled, serving as a proxy inspired by DOTS (sun2025improving), which evaluates the long-term effectiveness of relying almost entirely on implicit evidence without corrective feedback. Implementation details and further discussion of these baselines are provided in Appendix B.6.
@@
(1) BOTS achieves the best overall performance. For the 1.5B model, BOTS obtains 8 first-place and 9 second-place finishes out of 18 reported metrics, with a notable 36% acceleration (TTB(100%) = 0.64) in the math domain. For the 7B model, BOTS secures 6 first-place and 8 second-place finishes, including a remarkable 50% acceleration (TTB(100%) = 0.50) in the logic domain.
@@ 4.5 Overview of Additional Experiments
We provide extended experiments in Appendix F for a deeper understanding of BOTS: (1) A wall-clock breakdown (Appendix F.1) shows task selection adds less than 0.2% overhead. (2) Offline evaluation of the interpolator (Appendix F.2) confirms it effectively tracks task difficulty. (3) An ablation on Thompson sampling (Appendix F.3) shows posterior sampling yields more stable selection. (4) A fine-grained analysis of selected task dynamics (Appendix F.4) illustrates how BOTS shifts computation away from trivial ($p=1$) and impossible ($p=0$) tasks toward the informative mid-difficulty region.
@@ 5 Conclusion and Discussion
This work opens several promising directions. First, we focused mainly on binary-reward tasks; extending and validating BOTS in non-binary reward settings is an important next step. Second, BOTS currently uses fixed update rules, though our results show that different values of $\lambda$ and $\rho$ work best at different training stages. Developing adaptive update strategies that adjust to training dynamics would further improve robustness. Third, while our lightweight interpolator efficiently provides implicit evidence, designing stronger plug-in alternatives and systematically studying the trade-off between predictive accuracy and computational cost remain open challenges. A more detailed discussion of these directions is provided in Appendix E.4.
@@ Appendix A Notation Summary
$\lambda$ Forgetting factor in $[0,1]$ for historical information.
@@
Appendix B Implementational Details
B.1 Training Data
The training data for our experiments is sourced from GURU (cheng2025revisiting), a well-curated, cross-domain RL dataset covering mathematics, code, logic, science, simulation, and tabular tasks. Each subset has been rigorously deduplicated, verified, and filtered. Our training utilizes the math, code, and logic subsets, with the Zebra Puzzle excluded from the logic portion due to its non-binary reward structure.
@@ B.1 Training Data
The Code subset includes programming challenges from LeetCode (xia2025leetcodedataset), TACO-verified (likaixin2024taco-verified), PrimeIntellect (2025synthetic1), and LiveCodeBench (jain2024livecodebench). For PrimeIntellect and LiveCodeBench, it incorporates pre-filtered versions provided by DeepCoder [^2] https://www.together.ai/blog/deepcoder.
@@
The Logic subset is composed of several benchmarks designed to test structured reasoning. It includes Ordering Puzzles (relational ordering), Graph Puzzles (implicit graph traversal), the public training splits of ARC-AGI (chollet2024arc) and ARC-AGI-2 (chollet2025arc) (abstract grid transformations), and a 3.4k-sample from BARC (li2024barc) (synthetic ARC-style tasks). For these tasks, predictions are extracted from <answer> tags for reward calculation via exact match.
B.2 Evaluation Benchmarks
Followed by GURU (cheng2025revisiting), the evaluation suite consists of a set of established benchmarks to rigorously assess the model’s performance in a zero-shot setting across the same domains.
Math We evaluate mathematical reasoning on AIME24 (aime2024) and MATH500 (hendrycks2021measuringmathematicalproblemsolving), which cover a wide range of competition-level math problems.
Code Programming capabilities are assessed using HumanEval (chen202humaneval), MBPP (austin2021mbpp), and a subset of LiveCodeBench (jain2024livecodebench). These benchmarks span from basic function generation to complex algorithmic challenges.
Logic Logical and abstract reasoning are measured using Ordering Puzzles for general reasoning, and ARC-AGI (chollet2024arc) for grid-based abstract reasoning.
B.3 Training Details
The RFT setup follows standard practices (cheng2025revisiting; qu2025can; sun2025improving), employing GRPO (grpo2024) as the RL algorithm (detailed in Appendix B.5). All experiments are conducted with 8 NVIDIA A100 (80GB) GPUs. We utilize PyTorch’s FSDP for distributed training, with vLLM (kwon2023efficient) employed to accelerate the response generation during the rollout phase. The actor model is optimized using a learning rate of $1\times 10^{-6}$ and weight decay of 0.1. We apply a constant learning rate warmup for the first 10 steps and use gradient clipping with a maximum norm of 1.0. For the GRPO algorithm, we set the clipping ratio $\epsilon$ to 0.2 and generate 16 rollouts with a temperature of 1.0. The maximum prompt and response lengths are set to 4,096 and 8,192 tokens, respectively.
The training configurations are tailored for two LLMs: Qwen2.5-1.5B-Instruct and Qwen2.5-7B (Yang2024Qwen25TR), across three reasoning domains: Math, Code, and Logic. The primary differences across settings lie in batch sizes and memory optimization strategies to accommodate different model sizes. Specifically, the 7B model experiments utilize a larger training batch size (512 vs. 256) and GRPO mini-batch size (64 vs. 32). To fit the larger model on the same hardware, we enable FSDP’s CPU offloading for all 7B model experiments. Key hyperparameters for all six experimental settings are summarized in Table 6.

### v3

@@ 4.2 Analyzing the Impact of $\rho$
(2) Over-reliance on implicit evidence degrades long-term performance. As training progresses, settings with a large $\rho$ (e.g., 0.5, 1.0) show a declining ETR. This is driven by an increased tendency to select tasks whose success probability is $1$, as shown in the curve of <1<1 ratio. Consequently, the performance is negatively affected, as seen in the aggregated TTB and BSF metrics where large $\rho$ values underperform.
@@
4.3 Analyzing the Impact of $\lambda$
@@
In Section 3.5, we analyzed the role of the hyperparameter $\rho$ and its influence on BOTS. In this section, we empirically validate these analyses. We investigate the role of $\lambda$ in task selection by varying $\lambda\in\{0.0,0.05,0.1,0.2,0.5,1.0\}$, while keeping other hyperparameters fixed at their default values ($\rho=0.1$, posterior sampling enabled).
@@
Figure 3: Qwen2.5-1.5B-Instruct on Math with $\lambda\in\{0.0,0.05,0.1,0.2,0.5,1.0\},\rho=0.1$. Ratio of sampled training tasks (measured over 16 rollouts) with passing rates: strictly between 0 and 1 (L1), strictly greater than 0 (L2), and strictly less than 1 (L3), along with aggregated performance (MATH500, AMC23, and AIME24), plotted against training steps. All results are averaged over 3 random seeds, and 95% confidence intervals are reported. More detailed results are placed in Appendix G.7.
Benchmark MATH500 AMC23 AIME24 Math Aggregated Performance
Target Fraction 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100%
$\lambda=0.0$ 0.86 0.69 - 1.06 1.07 0.97 0.91 0.64 0.60 1.06 1.19 1.01 0.77 0.55 0.54 1.43 1.17 1.27 0.86 0.66 - 1.07 1.11 0.98
$\lambda=0.05$ 0.86 0.67 0.96 1.05 1.04 1.00 0.82 0.74 1.04 1.02 1.11 1.01 1.04 0.67 0.53 0.95 1.23 1.26 0.83 0.70 0.89 1.02 1.07 1.02
$\lambda=0.1$ 0.86 0.66 0.78 1.05 1.09 1.06 0.86 0.68 0.74 1.10 1.16 1.06 0.86 0.85 0.50 1.20 1.25 1.23 0.85 0.66 0.72 1.06 1.12 1.05
$\lambda=0.2$ 0.83 0.76 0.92 1.05 1.07 1.01 0.90 0.69 0.66 1.04 1.13 1.05 0.83 0.89 0.47 1.27 1.28 1.21 0.83 0.71 0.86 1.05 1.10 1.04
$\lambda=0.5$ 0.93 0.84 0.97 0.97 1.04 1.01 1.01 1.01 0.81 1.06 0.99 1.06 0.56 0.90 0.60 1.05 1.07 1.21 0.93 0.88 0.78 0.99 1.03 1.03
$\lambda=1.0$ 0.91 0.81 - 0.99 1.02 0.97 1.18 0.96 - 1.01 0.99 1.00 0.91 1.18 0.79 1.10 0.88 1.08 0.96 0.96 - 1.00 1.02 0.99
@@
Table 2: Qwen2.5-1.5B-Instruct on Math with $\lambda\in\{0.0,0.05,0.1,0.2,0.5,1.0\},\rho=0.1$. For TTB, notation “-” indicates that the target performance is never achieved within the evaluation window. The best and second best results are marked accordingly.
@@
Analysis and Takeaways. The hyperparameter $\lambda$ governs (i) the adaptivity of difficulty estimation under a non-stationary learning process, and (ii) the exploration–exploitation balance during task selection. Our key takeaway is that both of these aspects are crucial for strong long-term performance. An overly small $\lambda$ prevents the system from adapting to the model’s improving capability, whereas an overly large $\lambda$ injects excessive randomness into task selection. This conclusion is supported by the following observations from Figure 3 and Table 2:
@@
(1) Small $\lambda$ fails to track capability improvement. Settings with small $\lambda$ (e.g., 0.0, 0.05) exhibit a declining, especially <1<1 Ratio, in the mid-to-late training phase. This is because their preference to historical evaluations makes them slow to update the difficulty of tasks, causing ineffective tasks to be wastefully selected. This ultimately limits late-stage performance gains.
(2) Large $\lambda$ induces near-random task selection. Conversely, large $\lambda$ produce task selection behavior that is almost fully random. This arises because large $\lambda$ aggressively discounts historical evidence, yielding very small effective sample sizes and thus high posterior uncertainty, which in turn drives overly exploratory Thompson sampling behavior.
@@
(3) Moderate $\lambda$ achieves the best trade-off. A moderate $\lambda\in\{0.05,0.1,0.2\}$ strikes a balance: the difficulty estimates remain sufficiently adaptive while the uncertainty stays controlled, enabling a healthy exploration–exploitation balance. This leads to strong and consistent performance gains across the entire training process, achieving the best TTB and BSF results.
@@ 4.4 Performance Comparison Across Models and Domains
Offline 0.77 0.66 0.85 1.09 1.11 1.04 0.68 1.03 - 1.09 1.00 0.99 1.23 1.36 - 0.67 0.78 0.89
BOTS-MoPPS 1.02 0.90 0.89 0.98 1.03 1.00 0.78 0.96 1.09 1.09 1.02 1.02 0.87 1.04 - 1.13 0.96 0.96
BOTS-DOTS 0.70 0.70 0.73 1.08 1.08 1.01 0.67 0.65 0.79 1.17 1.09 1.02 0.66 1.32 - 1.28 0.92 0.93
BOTS 0.85 0.66 0.72 1.06 1.12 1.05 0.58 0.90 0.77 1.17 1.08 1.03 0.85 0.94 1.05 1.19 1.01 1.00
@@
We conduct extended experiments across model sizes (1.5B and 7B) and task domains (math, code, logic). We compare BOTS under the default setting ($\lambda=0.1,\rho=0.1$, posterior sampling enabled) against four baselines. Two are out-of-framework baselines: (i) Random, where tasks are uniformly sampled, and (ii) Offline, where tasks are ranked from easy to hard based on success probabilities of Qwen2.5-7B-Instruct and Qwen3-30B-A3B (for tie-breaking). Two are within-framework baselines: (iii) BOTS-MoPPS, with $\lambda=0.0,\rho=0.0$ and posterior sampling enabled, which reduces our framework to MoPPS (Qu et al., 2025) and thus enables direct evaluation of a purely explicit-evidence strategy; and (iv) BOTS-DOTS, with $\lambda=1.0,\rho=1.0$ and posterior sampling disabled, serving as a proxy inspired by DOTS (Sun et al., 2025), which evaluates the long-term effectiveness of relying almost entirely on implicit evidence without corrective feedback. Implementation details and further discussion of these baselines are provided in Appendix D.6.
@@
(1) BOTS achieves the best overall performance. For the 1.5B model, BOTS obtains 10 first-place and 6 second-place finishes out of 18 reported metrics, with a notable 28% acceleration (TTB(100%) = 0.72) in the math domain. For the 7B model, BOTS secures 6 first-place and 8 second-place finishes, including a remarkable 50% acceleration (TTB(100%) = 0.50) in the logic domain.
@@ 4.5 Overview of Additional Experiments
[…]
@@ 5 Conclusion and Discussion
This work opens several promising directions. First, we focused mainly on binary-reward tasks; extending and validating BOTS in non-binary reward settings is an important next step. Second, BOTS currently uses fixed update rules, though our results show that different values of $\lambda$ and $\rho$ work best at different training stages. Developing adaptive update strategies that adjust to training dynamics would further improve robustness. Third, while our lightweight interpolator efficiently provides implicit evidence, designing stronger plug-in alternatives and systematically studying the trade-off between predictive accuracy and computational cost remain open challenges. A more detailed discussion of these directions is provided in Appendix F.4.
Ethics Statement
All authors have read and adhered to the ICLR 2026 Code of Ethics. Our research focuses on the algorithmic efficiency of reinforcement finetuning for Large Language Models and does not involve human subjects, animal experiments, or the processing of personally identifiable information. The datasets used in our experiments are publicly available and established benchmarks within the research community; all software, datasets, and frameworks utilized are governed by the permissive Apache-2.0 open-source license. Our method aims to make AI research more sustainable and accessible, and we do not foresee any direct negative societal impacts or ethical concerns arising from our proposed methodology. The authors declare no conflict of interest.
Reproducibility Statement
We are committed to ensuring the full reproducibility of our research. To facilitate this, we will release our source code, which includes the implementation of the BOTS framework and the scripts required to replicate all experiments presented in this paper. The appendix contains a comprehensive description of our experimental setup, detailing all model configurations, dataset processing steps, and hyperparameter settings.
@@ Appendix A Notation Summary
$\lambda$ Discounting factor in $[0,1]$ for historical information.
@@
Appendix B Proofs
B.1 Proof for Proposition 3.2
[…]
Proof.
B.2 Proof for Proposition 3.5
[…]
Proof.
We drop the task index for clarity. From Equation (1) and Equation (3),
Appendix C Pseudo Codes
Algorithm 1 BOTS: Bayesian Online Task Selection
[…]
2:Trained model parameters $\theta$
3:Initialize $\bm{\alpha}_{0}$, $\bm{\beta}_{0}$. $\triangleright$ If not specified, $\bm{\alpha}_{0}=\mathbf{1}$, $\bm{\beta}_{0}=\mathbf{1}$
4:for $t=0$ to T1T-1 do
5:  if Thompson Sampling then
6:   Sample $\tilde{p}_{t}^{k}\sim\text{Beta}(\alpha_{t}^{k},\beta_{t}^{k}),k=1,\dots,N$
7:  else
8:   $\tilde{p}_{t}^{k}=\frac{\alpha_{t}^{k}}{\alpha_{t}^{k}+\beta_{t}^{k}},k=1,\dots,N$
9:  end if
10:  Select a batch of tasks $\{\mathcal{T}_{\mathcal{B}_{t}[i]}\}_{i=1}^{B}$ with top-$B$ utilities $\{\hat{u}_{k}:=-|\hat{p}_{k}-p^{*}|\}$.
11:  RFT with the selected tasks: $\theta_{t+1}=RFT(\theta_{t},\{\mathcal{T}_{\mathcal{B}_{t}[i]}\}_{i=1}^{B})$.
12:  Collect the online evaluation results of the selected tasks $\{r_{1:n}^{\mathcal{B}_{t}[i]}\}_{i=1}^{B}$.
13:  Compute explicit evidence $\{(s_{t}^{k},f_{t}^{k})\}_{k=1}^{N}$, with $\{r_{1:n}^{\mathcal{B}_{t}[i]}\}_{i=1}^{B}$. $\triangleright$ Eq. 2
14:  Compute implicit evidence $\{(\tilde{s}_{t}^{k},\tilde{f}_{t}^{k})\}_{k=1}^{N}$, with $\tilde{p}$, $\{r_{1:n}^{\mathcal{B}_{t}[i]}\}_{i=1}^{B}$. $\triangleright$ Eq. 3
15:  Update posterior $\bm{\alpha}_{t+1}$, $\bm{\beta}_{t+1}$, with $\lambda,\rho$. $\triangleright$ Eq. 1
16:end for
17:return $\theta_{T}$
Algorithm 2 Interpolation-based Implicit Evidence Estimator
1:Task index $k$, task batch $\{\mathcal{T}_{\mathcal{B}[i]}\}_{i=1}^{B}$, online evaluation results $\{r_{1:n}^{\mathcal{B}[i]}\}_{i=1}^{B}$, momentum coefficient $\gamma$.
2:Estimated success rate $\tilde{p}(k,\{\mathcal{T}_{\mathcal{B}[i]},r_{1:n}^{\mathcal{B}[i]}\}_{i=1}^{B})$ for $\mathcal{T}_{k}$.
3:Retrieve success rates of weak/strong reference models $\{(\bar{p}_{w}^{\mathcal{B}[i]},\bar{p}_{s}^{\mathcal{B}[i]})\}_{i=1}^{B}$. $\triangleright$ Pre-computed.
4:Compute the average empirical success rates of the current, weak, and strong models on $\mathcal{B}_{t}$ as
$\bar{p}^{\mathrm{ref}}(\mathcal{B})=\frac{1}{|\mathcal{B}|}\sum_{i=1}^{B}\frac{1}{n}\sum_{j=1}^{n}r_{j}^{\mathcal{B}[i]},\quad\bar{p}_{w}^{\mathrm{ref}}(\mathcal{B})=\frac{1}{|\mathcal{B}|}\sum_{i=1}^{B}\bar{p}_{w}^{\mathcal{B}[i]},\quad\bar{p}_{s}^{\mathrm{ref}}(\mathcal{B})=\frac{1}{|\mathcal{B}|}\sum_{i=1}^{B}\bar{p}_{s}^{\mathcal{B}[i]}.$
5:Estimate the relative capability coefficient of the current model as
$\mu(\mathcal{B})=\big(\bar{p}^{\mathrm{ref}}(\mathcal{B})-\bar{p}_{w}^{\mathrm{ref}}(\mathcal{B})\big)/\ \big(\bar{p}_{s}^{\mathrm{ref}}(\mathcal{B})-\bar{p}_{w}^{\mathrm{ref}}(\mathcal{B})\big).$
6:Momentum estimation: $\tilde{\mu}\leftarrow\gamma\tilde{\mu}+(1-\gamma)\mu(\mathcal{B})$.
7:Retirve success rates of weak/strong reference models on $\mathcal{T}_{k}$: $(\bar{p}_{w}^{k},\bar{p}_{s}^{k})$.
8:Linear interpolation with clipping: $\tilde{p}(k,\{\mathcal{T}_{\mathcal{B}[i]},r_{1:n}^{\mathcal{B}[i]}\}_{i=1}^{B})=\mathrm{clip}\,\Big(\tilde{\mu}\,\bar{p}_{s}^{k}+\bigl(1-\tilde{\mu}\bigr)\,\bar{p}_{w}^{k},\;0,\;1\Big).$
9:return $\tilde{p}(k,\{\mathcal{T}_{\mathcal{B}[i]},r_{1:n}^{\mathcal{B}[i]}\}_{i=1}^{B})$
Appendix D Implementational Details
D.1 Training Data
The training data for our experiments is sourced from GURU (Cheng et al., 2025), a well-curated, cross-domain RL dataset covering mathematics, code, logic, science, simulation, and tabular tasks. Each subset has been rigorously deduplicated, verified, and filtered. Our training utilizes the math, code, and logic subsets, with the Zebra Puzzle excluded from the logic portion due to its non-binary reward structure.
@@ D.1 Training Data
The Code subset includes programming challenges from LeetCode (Xia et al., 2025), TACO-verified (Li, 2024), PrimeIntellect (Mattern et al., 2025), and LiveCodeBench (Jain et al., 2024). For PrimeIntellect and LiveCodeBench, it incorporates pre-filtered versions provided by DeepCoder [^2] https://www.together.ai/blog/deepcoder.
@@
The Logic subset is composed of several benchmarks designed to test structured reasoning. It includes Ordering Puzzles (relational ordering), Graph Puzzles (implicit graph traversal), the public training splits of ARC-AGI (Chollet et al., 2024) and ARC-AGI-2 (Chollet et al., 2025) (abstract grid transformations), and a 3.4k-sample from BARC (Li et al., 2024b) (synthetic ARC-style tasks). For these tasks, predictions are extracted from <answer> tags for reward calculation via exact match.
D.2 Evaluation Benchmarks
Followed by GURU (Cheng et al., 2025), the evaluation suite consists of a set of established benchmarks to rigorously assess the model’s performance in a zero-shot setting across the same domains.
Math We evaluate mathematical reasoning on AIME24 (MAA, 2024) and MATH500 (Hendrycks et al., 2021), which cover a wide range of competition-level math problems.
Code Programming capabilities are assessed using HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and a subset of LiveCodeBench (Jain et al., 2024). These benchmarks span from basic function generation to complex algorithmic challenges.
Logic Logical and abstract reasoning are measured using Ordering Puzzles for general reasoning, and ARC-AGI (Chollet et al., 2024) for grid-based abstract reasoning.
D.3 Trainining Details
The RFT setup follows standard practices (Cheng et al., 2025; Qu et al., 2025; Sun et al., 2025), employing GRPO (Shao et al., 2024) as the RL algorithm (detailed in Appendix D.5). All experiments are conducted on verl framework (Sheng et al., 2025) with 8 NVIDIA A100 (80GB) GPUs. We utilize PyTorch’s FSDP for distributed training, with vLLM (Kwon et al., 2023) employed to accelerate the response generation during the rollout phase. The actor model is optimized using a learning rate of $1\times 10^{-6}$ and weight decay of 0.1. We apply a constant learning rate warmup for the first 10 steps and use gradient clipping with a maximum norm of 1.0. For the GRPO algorithm, we set the clipping ratio $\epsilon$ to 0.2 and generate 16 rollouts with a temperature of 1.0. The maximum prompt and response lengths are set to 4,096 and 8,192 tokens, respectively.
The training configurations are tailored for two LLMs: Qwen2.5-1.5B-Instruct and Qwen2.5-7B (Yang et al., 2024), across three reasoning domains: Math, Code, and Logic. The primary differences across settings lie in batch sizes and memory optimization strategies to accommodate different model sizes. Specifically, the 7B model experiments utilize a larger training batch size (512 vs. 256) and GRPO mini-batch size (64 vs. 32). To fit the larger model on the same hardware, we enable FSDP’s CPU offloading for all 7B model experiments. Key hyperparameters for all six experimental settings are summarized in Table 6.